OpenStudy (anonymous):

Could somebody help me figure this out? I keep having troubles with it. solve for x and y using substitution. 2x+4y=10 x=y-7

pooja195 (pooja195):

2(y-7)+4y=10 solve it

OpenStudy (anonymous):

2x+4y = 10 X= y-7 Substitute x into the first equation,.. So 2(y-7) + 4y= 10 2y-14 +4y=10 6y = 24 Y=4 Substitute for y into x=y-7 X=4-7 ** So x= -3 & y = 4 Hope this helps!